OpenStudy (anonymous):

solve 3/5x=7 1/2 what number is x

OpenStudy (anonymous):

please help

OpenStudy (maheshmeghwal9):

Is this the question? \[\frac{3}{5x}= 7\frac{1}{2}.\] ??????????????????

OpenStudy (anonymous):

yeah

OpenStudy (maheshmeghwal9):

\[\frac{3}{5x}=\frac{15}{2}.\] implies \[6=35x.\]

OpenStudy (maheshmeghwal9):

we obtained it after cross multiply.

OpenStudy (maheshmeghwal9):

so x = ?

OpenStudy (aravindg):

6=75x x=6/75

OpenStudy (anonymous):

well 6 doesnt show up on my quiz so im not sure

OpenStudy (maheshmeghwal9):

oops sorry 75x

OpenStudy (maheshmeghwal9):

@AravindG is right.

OpenStudy (anonymous):

it only show mixed fractions on my paper ?

OpenStudy (maheshmeghwal9):